OpenStudy (romantic_ch3micals):

HELP. Am I right? Medal. 26. Do the data in the table represent a direct variation or an inverse variation? write equation to model the data in the table. x: 1, 3, 5, 10 y: 4, 12, 20, 40 A. direct variation; y =4x B. direct variation; xy =1/4 C. inverse variation; xy =4 **** D. inverse variation; xy =1/4

OpenStudy (romantic_ch3micals):

@sleepyhead314 help?

OpenStudy (sleepyhead314):

you're saying that 10 times 40 equals 4 ?

OpenStudy (romantic_ch3micals):

oh...nevermind. It's A. I got confused.

OpenStudy (sleepyhead314):

there you go :P

OpenStudy (romantic_ch3micals):

Can you help me with one more. I'm not sure if it's correct.

OpenStudy (sleepyhead314):

of course

OpenStudy (romantic_ch3micals):

24. The temperature, T, of a given mass of gas varies inversely with its volume, V. The temperature of 20cm^3 of a certain gas is 15 degrees C. What will the temperature be when it is compressed to a volume of 5cm^3. A. 10C *** B. 300C C. 6.67C D. 60C

OpenStudy (sleepyhead314):

(20)*(15) = k k = ? then x = k / 5 = ?

OpenStudy (romantic_ch3micals):

D?

OpenStudy (romantic_ch3micals):

Because you divide 300 by 5...so I got 60

OpenStudy (sleepyhead314):

that's what I got ^_^

OpenStudy (romantic_ch3micals):

alright! I'll take my chances! Thanks

OpenStudy (sleepyhead314):

meadle?

OpenStudy (romantic_ch3micals):

Heh, sorry. I tend to forget